OpenStudy (anonymous):

Find the product of (3x - 7y)2.

OpenStudy (jojokiw3):

Distributive property. \[a(b + c) = ab + ac.\] Can you use this to figure it out?

OpenStudy (igreen):

\((3x - 7y)^2 \rightarrow (3x - 7y)(3x - 7y)\) Distribute by multiplying all the terms in the first parenthesis to all the terms in the 2nd parenthesis. 3x * 3x = ? 3x * -7y = ? -7y * 3x = ? -7y * -7y = ?

OpenStudy (anonymous):

Yes, it exactly

OpenStudy (jojokiw3):

Is she squaring it? Nevermind then. o_O